Chain Reaction, Inc., has been growing at a phenomenal rate of 30 percent per year because of its rapid expansion and explosive sales. You believe that this growth rate will last for three more years and that the rate will then drop to 10 percent per year. If the growth rate then remains at 10 percent indefinitely, what is the total value of the stock? Total dividend just paid were $5 million, and the required return is 20 percent. If there are 20 million shares, what is the value of stock per share?
Answer #1

Market Value = 5((1.3)/(1.2)) + 5((1.3)/(1.2))2 + 5((1.3)/(1.2))3 + 5(1.3)3(1.1)/(0.2 - 0.1)(1.2)3

Market Value = 5.42 + 5.86 + 6.36 + 69.93

Market Value = $87.57 million

Value of Stock per share = 87.57/20

Value of Stock per share = $4.3785